Copenhagen/11 August 2011



The A2 coloies we plated out yesterday, to check if they were coloured (no insert), were all white, so hopefully the have the insert.

Our instructor Kenneth is back. Eventhough the absoption spectras from yesterday on A1 didn't give a clear peak at 450 nm the protein can still be active, according to Kenneth. Today we will therefor give the protein substrate and hope that it will produce oximes.


  • Pcr on the A2 colonies and on an old A1, to check for insert. Both of them are in PSB1C3
  • TLC on A1 from the membrane preparation from yesterday
  • Try the user-cloning assembly from the DTU-team-2 with this protocol:USER-cloning

Back to Notebook